# Equivariant Graph Neural Operator for Modeling 3D Dynamics

### v1

@@ Abstract
Modeling the complex three-dimensional (3D) dynamics of relational systems is an important problem in the natural sciences, with applications ranging from molecular simulations to particle mechanics. Machine learning methods have achieved good success by learning graph neural networks to model spatial interactions. However, these approaches do not faithfully capture temporal correlations since they only model next-step predictions. In this work, we propose Equivariant Graph Neural Operator (EGNO), a novel and principled method that directly models dynamics as trajectories instead of just next-step prediction. Different from existing methods, EGNO explicitly learns the temporal evolution of 3D dynamics where we formulate the dynamics as a function over time and learn neural operators to approximate it. To capture the temporal correlations while keeping the intrinsic SE(3)-equivariance, we develop equivariant temporal convolutions parameterized in the Fourier space and build EGNO by stacking the Fourier layers over equivariant networks. EGNO is the first operator learning framework that is capable of modeling solution dynamics functions over time while retaining 3D equivariance. Comprehensive experiments in multiple domains, including particle simulations, human motion capture, and molecular dynamics, demonstrate the significantly superior performance of EGNO against existing methods, thanks to the equivariant temporal modeling.
@@
[…]
Equal contribution. Work done during Minkai’s internship at NVIDIA.
@@ 1 Introduction
Modeling the dynamics of many body systems in Euclidean space is an important problem in machine learning (ML) at all scales (Tenenbaum et al., 2011; Battaglia et al., 2016), ranging from interactions of atoms in a molecule or protein to interactions of celestial bodies in the universe. Traditional methods tackle these physical dynamics by Newton’s laws along with interactions calculated from physical rules, e.g., Coulomb force for modeling systems with charged particles (Kipf et al., 2018). Such calculations based on physical rules typically are very computationally expensive and slow, and recently researchers are studying alternative ML approaches by learning neural networks (NNs) to automatically capture the physical rules and model the interactions in a data-driven fashion. Along this research direction, considerable progress has been achieved in recent years by regarding objects as nodes and interactions as edges and learning the spatial interactions via graphs neural networks (GNNs). (Battaglia et al., 2016; Kipf et al., 2018; Sanchez-Gonzalez et al., 2019; Martinkus et al., 2021; Sanchez-Gonzalez et al., 2020; Pfaff et al., 2020).
Among the advancements, Equivariant GNN (EGNN) is a state-of-the-art approach with impressive results in modeling physical dynamics in 3D Euclidean space (Thomas et al., 2018; Satorras et al., 2021; Huang et al., 2022). EGNNs possess equivariance to roto-translational transformations in the Euclidean space, which has been demonstrated as a vital inductive bias to improve generalization (Köhler et al., 2019; Fuchs et al., 2020; Han et al., 2022; Xu et al., 2022). However, these methods typically model the dynamics by only learning to predict the next state given the current state, failing to faithfully capture the temporal correlation along the trajectory. As a result, performance is still often unsatisfactory without fully understanding the dynamics as systems evolve through time.
@@ 2.1 Graph Neural Networks
Graph neural networks (Gilmer et al., 2017; Kipf & Welling, 2017) are a family of neural network architectures for representation learning on relational structures, which has been widely adopted in modeling complex interactions and simulating physical dynamics. Early research in this field improves performance by designing expressive modules to capture the system interactions (Battaglia et al., 2016; Kipf et al., 2018; Mrowca et al., 2018) or imposing physical mechanics (Sanchez-Gonzalez et al., 2019). Beyond interactions, another crucial consideration for simulating physical dynamics is the symmetry in Euclidean space, i.e., the equivalence w.r.t rotations and translations. To this end, several works are first proposed to enforce translation equivariance (Ummenhofer et al., 2019; Sanchez-Gonzalez et al., 2020; Pfaff et al., 2020; Wu et al., 2018), and then later works further enforce the rotation equivariance by spherical harmonics (Thomas et al., 2018; Fuchs et al., 2020) or equivariant message passings (Satorras et al., 2021; Huang et al., 2022). In addition, some methods (Liu et al., 2022; Du et al., 2022) also introduce architectural enhancements with local frame construction to process geometric features while satisfying equivariance. However, despite the considerable progress, all the models are limited to modeling only the spatial interactions and ignore the critical temporal correlations, while in this paper, we view dynamics as a function of geometric states over time and explicitly learn the temporal correlations.
@@ 4.2 Equivariant Temporal Convolution Layer in Fourier Space
$\sigma({\mathcal{K}}_{\theta}f)=[\sigma({\mathcal{K}}_{\theta}f_{\mathbf{h}}),{\mathcal{K}}_{\theta}f_{\mathbf{Z}}]^{\textnormal{T}},$ (9)
which means we skip the activation layer for ${\mathbf{Z}}$ part and thus keep its original orientations. With the above designs, we enjoy the following crucial property:
@@ 5.2 Motion Capture
Dataset and implementation. We further benchmark our model on CMU Motion Capture dataset (CMU, 2003), which involves 3D trajectories captured from various human motion movements. We focus on two motions: Subject #35 (Walk) and Subject #9 (Run), following the setups and data splits in Huang et al. (2022); Han et al. (2022). Similar to N-body simulation, the input includes initial positions and velocities, and $\Delta T=30$. We use $P=5$ and uniform discretization by default. More details are in Appendix B.
@@ 5.3.2 Proteins
Dataset and implementation. We use the Adk equilibrium trajectory dataset (Seyler & Beckstein, 2017) integrated in the MDAnalysis (Richard J. Gowers et al., 2016) toolkit, which depicts the molecular dynamics trajectory of apo adenylate kinase. We follow the setting and split of Han et al. (2022). We use $P=4$ by default. We additionally consider EGHN (Han et al., 2022), the state-of-the-art model on this benchmark, as a baseline. To investigate the compatibility of our approach with different backbones, we incorporate equivariant temporal convolution into EGHN by inserting it to the end of each block in EGHN (details in Section B.4) and name it EGHNO.
@@ A.1 Statement of Universality
Line 43 follows since $\sigma$ is an identity operator on $\mathbf{Z}$ component. Lines 48 and 58 are the inverse and regular Fourier transforms and follows from Lemma A.2. The only new line is 53, which follows since $\mathbf{M}_{\theta}^{\mathbf{Z}}$ is a scalar multiplication for each component (which commutes with the matrix multiplication from $\mathbf{R}$). Therefore, we have that our neural operator is equivariant, as desired.
@@ B.1 Dataset Details
CMU Motion Capture. CMU Motion Capture dataset (CMU, 2003) involves 3D trajectories captured from various human motion movements. We focus on two motions: Subject #35 (Walk) and Subject #9 (Run), following the setups and data splits in Huang et al. (2022); Han et al. (2022). Subject #35 contains 200/600/600 trajectories for training/validation/testing, while Subject #9 contains 200/240/240. We view the joints as edges and their intersections (31 in total) as nodes. Similar to N-body simulation, the input includes initial positions and velocities of the intersections, and $\Delta T=30$. We use $P=5$ and uniform discretization by default.
Protein. We use the preprocessed version (Han et al., 2022) of the Adk equilibrium trajectory dataset (Seyler & Beckstein, 2017) integrated in the MDAnalysis (Richard J. Gowers et al., 2016) toolkit. In detail, the AdK equilibrium dataset depicts the molecular dynamics trajectory of apo adenylate kinase with CHARMM27 force field (MacKerell Jr et al., 2000), simulated with explicit water and ions in NPT at 300 K and 1 bar. The meta-data is saved every 240 ps for a total of 1.004 $\mu$s. We adopt the split by Han et al. (2022) which divides the entire trajectory into a training set with 2481 sub-trajectories, a validation set with 827, and a testing set with 878 trajectories, respectively. The backbone of the protein is extracted, and the graph is constructed using cutoff with radius 10Å.
@@ B.3 Hyperparameters
We provide detailed hyperparameters of our EGNO in Table 8. Specifically, batch is for batch size, lr for learning rate, wd for weight decay, layer for the number of layers, hidden for hidden dimension, timestep for the number of time steps, time_emb for the dimension of time embedding, num_mode for the number of modes (frequencies). We adopt Adam optimizer (Kingma & Ba, 2014) and all models are trained towards convergence with an earlystopping of 50 epochs on the validation loss. In particular, for EGHNO, we require additional hyperparemeters: the number of pooling layers is 4, the number of decoding layers is 2, and the number of message passing layer is 4, the same as Han et al. (2022). For the baselines, we strictly follow the setup in previous works and report the numbers from them in the S2S setting. For S2T, we keep the backbone EGNN the same number of layers and hidden dimension as ours for EGNN-R and EGNN-S, ensuring fairness.
@@ B.4 Model Details of EGHNO
Here we provide an illustration of EGHNO in Figure 4 to show the details of its architecture. In addition to the equivariant GNN layers, we have the equivariant pooling and unpooling layers introduced in the equivariant graph hierarchical network (EGHN) (Han et al., 2022). To realize EGHNO, we also stack the temporal convolution layers before each EGNN layer similar to EGNO (Figure 1). In particular, we follow EGHN and add jump connections between the equivariant unpooling layer and the corresponding low-level pooling layer, which helps to capture the low-level details of large geometric systems.

### v2

@@ Abstract
Modeling the complex three-dimensional (3D) dynamics of relational systems is an important problem in the natural sciences, with applications ranging from molecular simulations to particle mechanics. Machine learning methods have achieved good success by learning graph neural networks to model spatial interactions. However, these approaches do not faithfully capture temporal correlations since they only model next-step predictions. In this work, we propose Equivariant Graph Neural Operator (EGNO), a novel and principled method that directly models dynamics as trajectories instead of just next-step prediction. Different from existing methods, EGNO explicitly learns the temporal evolution of 3D dynamics where we formulate the dynamics as a function over time and learn neural operators to approximate it. To capture the temporal correlations while keeping the intrinsic SE(3)-equivariance, we develop equivariant temporal convolutions parameterized in the Fourier space and build EGNO by stacking the Fourier layers over equivariant networks. EGNO is the first operator learning framework that is capable of modeling solution dynamics functions over time while retaining 3D equivariance. Comprehensive experiments in multiple domains, including particle simulations, human motion capture, and molecular dynamics, demonstrate the significantly superior performance of EGNO against existing methods, thanks to the equivariant temporal modeling. Our code is available at https://github.com/MinkaiXu/egno.
@@ 1 Introduction
Modeling the dynamics of many body systems in Euclidean space is an important problem in machine learning (ML) at all scales (Tenenbaum et al., 2011; Battaglia et al., 2016), ranging from interactions of atoms in a molecule or protein to interactions of celestial bodies in the universe. Traditional methods tackle these physical dynamics by Newton’s laws along with interactions calculated from physical rules, e.g., Coulomb force for modeling systems with charged particles (Kipf et al., 2018). Such calculations based on physical rules typically are very computationally expensive and slow, and recently researchers are studying alternative ML approaches by learning neural networks (NNs) to automatically capture the physical rules and model the interactions in a data-driven fashion. Along this research direction, considerable progress has been achieved in recent years by regarding objects as nodes and interactions as edges and learning the spatial interactions via graphs neural networks (GNNs) (Battaglia et al., 2016; Kipf et al., 2018; Sanchez-Gonzalez et al., 2019; Martinkus et al., 2021; Sanchez-Gonzalez et al., 2020; Pfaff et al., 2020).
Among the advancements, Equivariant GNN (EGNN) is a state-of-the-art approach with impressive results in modeling physical dynamics in 3D Euclidean space (Thomas et al., 2018; Satorras et al., 2021). EGNNs possess equivariance to roto-translational transformations in the Euclidean space, which has been demonstrated as a vital inductive bias to improve generalization (Köhler et al., 2019; Fuchs et al., 2020; Han et al., 2022b; Xu et al., 2022). However, these methods typically model the dynamics by only learning to predict the next state given the current state, failing to faithfully capture the temporal correlation along the trajectory (Song et al., 2023; Song & Dhariwal, 2024). As a result, performance is still often unsatisfactory without fully understanding the dynamics as systems evolve through time.
@@ 2.1 Graph Neural Networks
Graph neural networks (Gilmer et al., 2017; Kipf & Welling, 2017) are a family of neural network architectures for representation learning on relational structures, which has been widely adopted in modeling complex interactions and simulating physical dynamics. Early research in this field improves performance by designing expressive modules to capture the system interactions (Battaglia et al., 2016; Kipf et al., 2018; Mrowca et al., 2018) or imposing physical mechanics (Sanchez-Gonzalez et al., 2019). Beyond interactions, another crucial consideration for simulating physical dynamics is the symmetry in Euclidean space, i.e., the equivalence w.r.t rotations and translations. To this end, several works are first proposed to enforce translation equivariance (Ummenhofer et al., 2019; Sanchez-Gonzalez et al., 2020; Pfaff et al., 2020; Wu et al., 2018), and then later works further enforce the rotation equivariance by spherical harmonics (Thomas et al., 2018; Fuchs et al., 2020) or equivariant message passings (Satorras et al., 2021; Huang et al., 2022; Han et al., 2022a). In addition, some methods (Liu et al., 2022; Du et al., 2022) also introduce architectural enhancements with local frame construction to process geometric features while satisfying equivariance. However, despite the considerable progress, all the models are limited to modeling only the spatial interactions and ignore the critical temporal correlations, while in this paper, we view dynamics as a function of geometric states over time and explicitly learn the temporal correlations.
@@ 4.2 Equivariant Temporal Convolution Layer in Fourier Space
$\sigma({\mathcal{K}}_{\theta}f)=[\sigma({\mathcal{K}}_{\theta}f_{\mathbf{h}}),\hat{\sigma}({\mathcal{K}}_{\theta}f_{\mathbf{Z}})]^{\textnormal{T}},$ (9)
where $\hat{\sigma}$ denotes equivariant nonlinear activation layers for ${\mathbf{Z}}$ part. It can be realized as $\hat{\sigma}({\mathcal{K}}_{\theta}f_{\mathbf{Z}})=\sigma(||{\mathcal{K}}_{\theta}f_{\mathbf{Z}}||)\cdot{\mathcal{K}}_{\theta}f_{\mathbf{Z}}$, which is a scalar transform of the directional feature and thus keep its original orientations (Thomas et al., 2018). We in practice did not observe significant performance change by introducing nonlinear onto the ${\mathbf{Z}}$. With the above designs, we enjoy the following crucial property:
@@ 5.2 Motion Capture
Dataset and implementation. We further benchmark our model on CMU Motion Capture dataset (CMU, 2003), which involves 3D trajectories captured from various human motion movements. We focus on two motions: Subject #35 (Walk) and Subject #9 (Run), following the setups and data splits in Huang et al. (2022); Han et al. (2022b). Similar to N-body simulation, the input includes initial positions and velocities, and $\Delta T=30$. We use $P=5$ and uniform discretization by default. More details are in Appendix B.
@@ 5.3.2 Proteins
Dataset and implementation. We use the Adk equilibrium trajectory dataset (Seyler & Beckstein, 2017) integrated in the MDAnalysis (Richard J. Gowers et al., 2016) toolkit, which depicts the molecular dynamics trajectory of apo adenylate kinase. We follow the setting and split of Han et al. (2022b). We use $P=4$ by default. We additionally consider EGHN (Han et al., 2022b), the state-of-the-art model on this benchmark, as a baseline. To investigate the compatibility of our approach with different backbones, we incorporate equivariant temporal convolution into EGHN by inserting it to the end of each block in EGHN (details in Section B.4) and name it EGHNO.
@@ 6 Conclusion
Acknowledgments
We thank Hongkai Zheng for constructive methodology suggestions, Vignesh Bhethanabotla and Shengchao Liu for general discussions. We gratefully acknowledge the support of HAI, NSF(#1651565), ARO (W911NF-21-1-0125), ONR (N00014-23-1-2159), CZ Biohub. We also gratefully acknowledge the support of DARPA under Nos. N660011924033 (MCS); NSF under Nos. OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Stanford Institute for Human-Centered AI, Chan Zuckerberg Initiative, Amazon, Genentech, GSK, Hitachi, SAP, and UCB. Minkai Xu thanks the generous support of Sequoia Capital Stanford Graduate Fellowship.
Impact Statement
This paper presents a temporal equivariant neural operator for modeling 3D dynamics, with the goal to advance fundamental Machine Learning research. There are some potential societal consequences of our work that are related to 3D dynamics modeling, none of which we feel must be specifically highlighted here.
@@ A.1 Statement of Universality
Line 22 follows since $\sigma$ is an identity operator on $\mathbf{Z}$ component. Lines 23 and 25 are the inverse and regular Fourier transforms and follows from Lemma A.2. The only new line is 24, which follows since $\mathbf{M}_{\theta}^{\mathbf{Z}}$ is a scalar multiplication for each component (which commutes with the matrix multiplication from $\mathbf{R}$). Therefore, we have that our neural operator is equivariant, as desired.
@@ B.1 Dataset Details
CMU Motion Capture. CMU Motion Capture dataset (CMU, 2003) involves 3D trajectories captured from various human motion movements. We focus on two motions: Subject #35 (Walk) and Subject #9 (Run), following the setups and data splits in Huang et al. (2022); Han et al. (2022b). Subject #35 contains 200/600/600 trajectories for training/validation/testing, while Subject #9 contains 200/240/240. We view the joints as edges and their intersections (31 in total) as nodes. Similar to N-body simulation, the input includes initial positions and velocities of the intersections, and $\Delta T=30$. We use $P=5$ and uniform discretization by default.
Protein. We use the preprocessed version (Han et al., 2022b) of the Adk equilibrium trajectory dataset (Seyler & Beckstein, 2017) integrated in the MDAnalysis (Richard J. Gowers et al., 2016) toolkit. In detail, the AdK equilibrium dataset depicts the molecular dynamics trajectory of apo adenylate kinase with CHARMM27 force field (MacKerell Jr et al., 2000), simulated with explicit water and ions in NPT at 300 K and 1 bar. The meta-data is saved every 240 ps for a total of 1.004 $\mu$s. We adopt the split by Han et al. (2022b) which divides the entire trajectory into a training set with 2481 sub-trajectories, a validation set with 827, and a testing set with 878 trajectories, respectively. The backbone of the protein is extracted, and the graph is constructed using cutoff with radius 10Å.
@@ B.3 Hyperparameters
We provide detailed hyperparameters of our EGNO in Table 8. Specifically, batch is for batch size, lr for learning rate, wd for weight decay, layer for the number of layers, hidden for hidden dimension, timestep for the number of time steps, time_emb for the dimension of time embedding, num_mode for the number of modes (frequencies). We adopt Adam optimizer (Kingma & Ba, 2014) and all models are trained towards convergence with an earlystopping of 50 epochs on the validation loss. In particular, for EGHNO, we require additional hyperparemeters: the number of pooling layers is 4, the number of decoding layers is 2, and the number of message passing layer is 4, the same as Han et al. (2022b). For the baselines, we strictly follow the setup in previous works and report the numbers from them in the S2S setting. For S2T, we keep the backbone EGNN the same number of layers and hidden dimension as ours for EGNN-R and EGNN-S, ensuring fairness.
@@ B.4 Model Details of EGHNO
Here we provide an illustration of EGHNO in Figure 4 to show the details of its architecture. In addition to the equivariant GNN layers, we have the equivariant pooling and unpooling layers introduced in the equivariant graph hierarchical network (EGHN) (Han et al., 2022b). To realize EGHNO, we also stack the temporal convolution layers before each EGNN layer similar to EGNO (Figure 1). In particular, we follow EGHN and add jump connections between the equivariant unpooling layer and the corresponding low-level pooling layer, which helps to capture the low-level details of large geometric systems.